COMPLETELY LOST HELP NEEDED ON HOMEWORK QUESTION IT would be much appreciated
Given the function f(x)=20sin(30*x)-10 determine when f(x)=-20

Answer by stanbon(75887)

f(x)=20sin(30*x)-10 determine when f(x)=-20
----
20sin(30x) - 10 = -20
---
20sin(30x) = -10
---
sin(30x) = -1/2
----
30x = -pi/6 or 30x = 7pi/6)
---
In radians:
x = -pi/180 or x = (7pi)/180
----------------------
In degrees:
x = -1 degree or x = 7 degrees
===============
Cheers,
Stan H.